no08189:

Point A is at (3, 2.6) and point B is at (3, 1.4) on a coordinate grid. The distance between the two points is ____. Input numbers and decimal point only, such as 5.8.

xxDeppressionxx:

Is this math

no08189:

yes

wolfkid:

yeah

xxDeppressionxx:

2.6 - 1.4 = 1.2

XioGonz:

Notice how the x in each point is 3 let's not touch that So we got to subtract. 2.6 - 1.4 = 1.2

xxDeppressionxx:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @XioGonz Notice how the x in each point is 3 let's not touch that So we got to subtract. 2.6 - 1.4 = 1.2 \(\color{#0cbb34}{\text{End of Quote}}\) Site your source

xxDeppressionxx:

the answer is 1.2

no08189:

ok thx

xxDeppressionxx:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @no08189 ok thx \(\color{#0cbb34}{\text{End of Quote}}\) Np

xxDeppressionxx:

Oh hyeah and Welcome to Qc

no08189:

thx

xxDeppressionxx:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @no08189 thx \(\color{#0cbb34}{\text{End of Quote}}\) Your Welcome